OpenStudy (anonymous):

Find all zeros in the equation. x^4+8x^2-9=0 Can you please explain what steps to take, so I can do every other question in my practice problems? Thank you.

OpenStudy (anonymous):

\[x^4+8x^2-9=0\] if it factors, use that.

OpenStudy (anonymous):

if it is does not factor you can replace x by z^2 and solve \[z^2+8z-9=0\] but i think this factors as \[(x^2+9)(x^2-1)=0\] therefore \[x^2+9=0\] or \[x^2-1=0\]

OpenStudy (anonymous):

if you are working over the real numbers, there is no solution to \[x^2+9=0\] because \[x^2\geq0\] and therefore it cannot be -9

OpenStudy (anonymous):

and then \[x^2-1=0\] \[(x+1)(x-1)=0\] \[x=1 \text { or } x = -1\] or you can write \[x^2-1=0\implies x^2=1\implies x=\pm1\]

OpenStudy (anonymous):

okay. thank you.

OpenStudy (anonymous):

yw